OpenStudy (acespeedfighter):

Part A: Explain why x = 3 makes 4x − 1 ≤ 11 true but not 4x − 1 < 11. (5 points) Part B: What value from the set {4, 5, 6, 7, 8} makes the equation 8x + 1 = 57 true? Show your work. (5 points)

OpenStudy (pythagoras123):

4(3)-1 = 11 11 = 11 Part B: 8x = 56 x = 7 (answer).

OpenStudy (acespeedfighter):

ty

OpenStudy (acespeedfighter):

I need it to be explained in word

OpenStudy (pythagoras123):

4x − 1 ≤ 11 This means that 4x-1 can be less than OR equal to 11. (note the usage of the word "or") When x = 4, 4x-1 = 11. Hence x=4 satisfies the inequality.

OpenStudy (pythagoras123):

Part B: Simply explain that x=7 is the only solution of the equation 8x + 1 = 57.

OpenStudy (acespeedfighter):

oh ok thx :D